OpenStudy (anonymous):

graph the exponential model y =3(6)^x i just need the points

OpenStudy (barrycarter):

What happens when you plug in x=1, x=2, and so on?

OpenStudy (cwrw238):

when x = 0 y = 3(6)^0 0 can you calculate that?

OpenStudy (cwrw238):

what is a number to power zero equal to?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

so it would be 1, 18

OpenStudy (cwrw238):

right so when x=0, y = 3*1 = 3 6^0 = 1

OpenStudy (anonymous):

thank you so so much :D

OpenStudy (cwrw238):

yw - the other points are straight forward it will be a steep curve something like this 

OpenStudy (anonymous):

ohhh i gotcha

OpenStudy (cwrw238):

good